# Unveiling Topological Structures in Text: A Comprehensive Survey of Topological Data Analysis Applications in NLP

### v2

@@ 4.2.1 TF-IDF
TF-IDF (Term Frequency - Inverse Document Frequency) is a well-known statistical formula that calculates the importance of words relative to a corpus. A few works investigated the extraction of topological features from TF-IDF representations as part of the pipeline illustrated in Fig. 4. For instance, SIFT, a persistent homology-based model with TF-IDF, is developed to differentiate between child and adolescent writings Zhu (2013). It is later adopted for deepfake text detection Løvlie (2023), extracting features from presidential election speeches from 2000-2020 Huang (2022), and movie genre classification with over 15% increase in F1 score Doshi and Zadrozny (2018); Shin (2019). Some other applications include - legal entailment Savle et al. (2019), sentiment analysis Michel et al. (2017), keyphrase extraction Guan et al. (2016), twitter topic detection Torres-Tramón et al. (2015), and text summarization Kumar and Sarkar (2022). In addition, TF-IDF has also been used as input for mapper to explain linguistic phenomena such as geographical differences of nursery rhymes Maadarani and Hajra (2020) and behaviors of NLP model van Veen (2020). Elyasi and Moghadam (2019) compares the two popular TDA approaches - Persistent Homology and Mapper to classify Persian poems. Lastly, Effah (2017) applies TDA to age group categorization of lonely people.
@@ 4.2.2 Pre-trained Non-contextual Embeddings
Word2Vec embeddings are a type of word representation that allows words with similar meanings to have similar vector representations Mikolov et al. (2013). Word2vec has been used as inputs to persistent homology to extract a hierarchical document structure, known as story trees, that allows tracing storylines Haghighatkhah et al. (2022), or with mapper to construct a graph of word2vec embeddings to create a topological search engine Cornell (2020). TDA has also been applied on Word2vec to solve innovative problems such as measuring the distance between the literary style of Spanish poets belonging to similar genres Paluzo Hidalgo et al. (2019), the detection of fraudulent papers Tymochko et al. (2021), detecting topological loops in logical statements Tymochko et al. (2020), disease prediction from epidemic curves Petri and Leitao (2020), comparing between embeddings of text and speech Yessenbayev and Kozhirbayev (2022, 2024). TDA features extracted from Word2vec have also been used for unsupervised learning tasks such as topic modeling Holmes (2020), word sense induction and disambiguation Rawson et al. (2022); Temčinas (2018), derivation of the correlation between sentence vectors and their semantics Sun and Nelson (2023) or distinguishing subsets of Wikipedia data Wright and Zheng (2020). Such features have also been compared with those extracted from GloVe or LLMs’ embeddings using persistent homology Feng et al. (2024).
@@
GloVe or Global Vectors for Word Representation is another technique for numerically representing texts as embeddings. GloVe has been used to extract useful topological features for classification (i.e., topological signatures of whole texts) such as those of novel writers Gholizadeh et al. (2018), time-series-based sliding windows Gholizadeh et al. (2020), and keywords Novak (2019). These features have also been used for unique applications such as comparing cross-lingual sentence representations Haim Meirom and Bobrowski (2022), explaining model performance Spannaus et al. (2024), social anxiety detection Byers (2021), and deepfake text detection Løvlie (2023). TDA features extracted from GloVe have also been compared with classical text embedding such as TF-IDF Michel et al. (2017) and also combined with those extracted from BERT and BiLSTM Deng and Duzhin (2022) in improving the performance of both classification and clustering tasks. Other applications include - extracting logic and reasoning in texts to test manifestation of intelligence and understanding in models Zadrozny (2021a), and capturing circles in circular arguments Tymochko et al. (2020); Zadrozny (2021b).
@@
FastText embeddings build upon the Word2Vec approach by incorporating subword information, enhancing the representation of rare words, and allowing for embedding out-of-vocabulary words Bojanowski et al. (2017). Jakubowski et al. (2020), Triki (2021), and Shehu (2024) use persistent homology to measure the polysemy of a word. They find that topological polysemy is inconsistent across different word embeddings. Jakubowski et al. (2020) further uses a similar technique to perform word sense induction & disambiguation from SemEval-2010 task. Also, Temčinas (2018) performs unsupervised word sense induction & disambiguation. Tymochko et al. (2021) extracts topological features from FastText embeddings to detect fraudulent papers.
@@ 4.2.3 Pre-trained Contextual Embeddings
Researchers evaluated the strength of the TDA features extracted from Transformer-based Vaswani et al. (2017) embeddings. Using the idea of self-attention, the neural network can encode more semantic and syntactic features than previous embeddings which should allow for richer TDA features to be extracted. To incorporate TDA features for classification, several researchers have investigated the efficacy of using other outputs of the encoder models - CLS Embedding output, hidden weights and attention weights to extract high-quality additional features.
CLS Embedding Output. Tulchinskii et al. (2024) proposes a Persistent Homology Dimension (PHD) estimator that measures the intrinsic dimensionality of the manifold underlying the set of CLS embeddings for a given text sample for deepfake text detection. PHD has also been utilized to perform boundary detection (i.e., where the human author stops and the AI author begins) in the context of deepfake text detection Kushnareva et al. (2024). CLS embeddings have been utilized to extract TDA features for other supervised learning tasks such as fake news detection Lavery et al. (2024), TEDtalk public speaking ratings classification Das et al. (2021) or unsupervised learning tasks such as topic modeling Byrne et al. (2022) and estimating class separability of text datasets Gourgoulias et al. (2024). Extracted topological features from CLS embeddings have also helped selecting the best ensemble model Proskura and Zaytsev (2024a) and visualizing the training process of transformer-based models Rathore et al. (2023).
@@
Hidden Weights. Uchendu et al. (2024) propose TopFormer, an end-to-end fine-tuned Topological RoBERTa-base model for deepfake text attribution by using the reshaped $pooled\_output$ (to 2D) as input for the TDA layer. Similar to Jakubowski et al. (2020), Triki (2021), and Shehu (2024), Garcia (2022) studies polysemous words. Using Mapper to visualize the hidden representations of polysemous words in the BERT transformer model, Garcia (2022) studies how BERT represents the semantics of ambiguous words. Similarly, Gardinazzi et al. (2024) proposes a novel metric - persistence similarity to describe the topological properties of internal representations in LLMs, detect topologically similar layers in an LLM, and prune redundant layers. Sun and Nelson (2023) derives the correlation between sentence vectors and their semantics using persistent homology. Chauhan and Kaul (2022) proposes a novel scoring metric - persistence scoring function which captures the homology of the hidden representations of BERT, correlates with test accuracy, and is robust to adversarial perturbations. García-Castellanos et al. (2024) performs zero-shot model stitching by employing topological densification to encourage clustering of similar points. In addition, Ruppik et al. (2024) extracts topological features using persistent homology from RoBERTa’s hidden weights for dialogue term extraction. Finally, Kudriashov et al. (2024) probes BERT’s hidden weights using persistent homology when training BERT on new grammatical features, known as polypersonality.
Attention Weights. BERT’s attention weights have been transformed to both directed and undirected graphs, on top of which different TDA features are extracted for deepfake text detection Kushnareva et al. (2021), investigation of whether language models learn grammatical concepts central to human linguistic competence (i.e., grammatical acceptability judgment) in different languages including English, Italian, Swedish, Russian Cherniavskii et al. (2022); Proskurina et al. (2023) and speech recognition task Tulchinskii et al. (2023). Similar techniques have also been used on RoBERTa’s attention weights for similar and other important tasks in linguistics such as grammatical checks Perez and Reinauer (2022), dialog term extraction Vukovic et al. (2022), deepfake text detection Uchendu et al. (2024), and document coherence Jain et al. (2024).
TDA features extracted from attention weights have also been used for model uncertainty estimation Kostenok et al. (2023), robustness evaluation of TDA features Perez and Reinauer (2022), and out-of-distribution detection (OOD) Pollano et al. (2024), dynamic weighting for ensemble models Proskura and Zaytsev (2024b). Especially,  Pollano et al. (2024) find that models that utilize TDA features extracted from attention weights perform best on far out-of-domain samples compared to those that use TDA features extracted from CLS outputs Pollano et al. (2024).
@@
ELMo Embeddings. ELMo embeddings are a type of word representation that captures both the meaning of words and their usage in context Peters et al. (2018). Similar to other embeddings, ELMo has also been leveraged to extract topological features. Tymochko et al. (2021) uses ELMo embeddings to extract topological features for detecting fraudulent papers by examining their titles and abstracts. To compare the strength of ELMo embeddings, they further extract topological features with other embeddings (Word2Vec, GloVe, FastText, and Frequency Time Series) to ascertain the best embeddings to extract strong topological features.
@@ 4.2.4 Symbolic Representations
PLC refers to rules and methods used to encode letters that fuel various communication systems, cryptography techniques, or linguistic analyses. Letter coding transforms letters or characters into different symbols, numbers, or other forms. PSSC is similar to PLC but for extracting topological features from speech sounds. One particular application of PLC and PSSC is the study of tongue twisters, such as for Ukrainian Yurchuk and Gurnik (2023); Kovaliuk et al. (2024).
@@ 4.2.5 Multi-Modal Representations
TDA features have also been extracted from other NLP-related feature representations, including multi-media data such as audio and video. There have been a variety of research utilizing such multi-modal inputs to TDA layers, ranging from studying human vowels and infant vocalizations Bonafos et al. (2023, 2024); Noakes (1991), emotion recognition from audio speech Gonzalez-Diaz et al. (2019) and videos Paluzo-Hidalgo et al. (2022), depression detection from audio clips Tlachac et al. (2020) to recognizing voiced and voiceless consonants in speech Zhu et al. (2024) and music classification Bergomi (2015); Sassone et al. (2022), the majority of which utilizes persistent homology to extract TDA features. Due to the multi-modality nature of these works, they also help inform novel linguistic-sound findings such as consonants can be detected using vibration patterns that are captured by persistent homology Zhu et al. (2024).

### v3

@@ 4.2.1 TF-IDF
TF-IDF (Term Frequency - Inverse Document Frequency) is a well-known statistical formula that calculates the importance of words relative to a corpus. A few works investigated the extraction of topological features from TF-IDF representations as part of the pipeline illustrated in Fig. 4. For instance, SIFT, a persistent homology-based model with TF-IDF, is developed to differentiate between child and adolescent writings Zhu (2013). This model represented the TF-IDF features as a time series, and then extracted topological features to enhance text classification. Several other researchers applied this model to other classification tasks, such as deepfake text detection Løvlie (2023), presidential election speech attribution Huang (2022), and movie genre classification Doshi and Zadrozny (2018); Shin (2019).
Insight (TFIDF). Topological features extracted from TF-IDF have only been applied to the text classification task. It is possible that topological features extracted from TF-IDF are not rich enough to perform the other six tasks.
@@ 4.2.2 Pre-trained Non-contextual Embeddings
Word2Vec embeddings are a type of word representation that allows words with similar meanings to have similar vector representations Mikolov et al. (2013). Unlike TF-IDF, topological features extracted from word2vec have been applied to several tasks, including structure & visualization, where Haghighatkhah et al. (2022) creates story trees to trace story lines. Next, we observe applications in sentiment & semantic analysis task, where TDA is applied to novel problems such as the creation of a topological search engine Cornell (2020), measuring distance between the literary style of Spanish poets Paluzo Hidalgo et al. (2019), word sense induction and disambiguation Rawson et al. (2022); Temčinas (2018), derivation of the correlation between sentence vectors and their semantics Sun and Nelson (2023), and comparing text and speech embeddings Yessenbayev and Kozhirbayev (2022, 2024). Thirdly, for the classification task, researchers detect fraudulent papers Tymochko et al. (2021), and topological loops in logical statements Tymochko et al. (2020).
Additionally, we observe applications in the health, social, and scholarly analysis task - disease prediction from epidemic curves Petri and Leitao (2020), topic modeling tasks Holmes (2020); Wright and Zheng (2020). Finally, for the model interpretation and analysis task, Feng et al. (2024) uses both topological and geometrical features to investigate the strength of LLM-enhanced data augmentation.
Insight (Word2Vec). Topological features extracted from Word2vec are considered rich, such that researchers have attempted six out of seven tasks. The only tasks that have not been attempted are speech and music processing, since different embedding is needed for audio data. Finally, we observe that the sentiment and semantic analysis task is the most popular task for word2vec.
@@
GloVe or Global Vectors for Word Representation is another technique for numerically representing texts as embeddings. The tasks we observe that are attempted using topological features extracted from GloVe embeddings are - classification, where subtasks include novel author attribution Gholizadeh et al. (2018) , deepfake text detection Løvlie (2023); sentiment and semantic analysis Gholizadeh et al. (2020); Tymochko et al. (2020); Zadrozny (2021b); model interpretation and analysis, where Haim Meirom and Bobrowski (2022); Michel et al. (2017) compare text representations and embeddings, Spannaus et al. (2024) explains model performance, and Zadrozny (2021a) tests the manifestation of intelligence and understanding in models; health, social, and scholarly analysis Byers (2021); Novak (2019); and clustering and topic modeling Deng and Duzhin (2022)
Insight (GloVe). Topological features extracted from GloVe are considered rich, such that researchers have attempted all tasks, except (1) speech and music processing and (2) structure and visualization. The most popular application is model interpretation and analysis, suggesting that glove embeddings can be sufficiently probed using TDA techniques to excavate explanations for performance.
@@
FastText embeddings build on the Word2Vec approach by incorporating subword information, improving the representation of rare words, and allowing for embedding out-of-vocabulary words Bojanowski et al. (2017). This type of embedding is not a popular feature researchers use to enhance topological features as only two tasks are attempted - sentiment and semantic analysis Jakubowski et al. (2020); Triki (2021); Shehu (2024), specifically for word sense induction & disambiguation; and text classification, where Tymochko et al. (2021) detects fraudulent papers.
Insight (FastText). FastText is not popularly adopted for topological applications, with only a few researchers applying it to two tasks - text classification and sentiment & semantic analysis.
@@ 4.2.3 Pre-trained Contextual Embeddings
Researchers evaluated the strength of the TDA features extracted from Transformer-based Vaswani et al. (2017) embeddings. Using the idea of self-attention, the neural network can encode more semantic and syntactic features than previous embeddings which should allow for richer TDA features to be extracted. To incorporate TDA features for various tasks, several researchers have investigated the efficacy of using other outputs of encoder and decoder models - CLS Embedding output, hidden weights, and attention weights to extract high-quality additional features.
CLS Embedding Output. Researchers have applied these features on text classification tasks, specifically for deepfake text detection Tulchinskii et al. (2024); Kushnareva et al. (2024); Wei et al. (2025); Guilinger et al. (2025), fake news detection Lavery et al. (2024), and TEDtalk public speaking ratings classification Das et al. (2021).
Additionally, we observe applications to the topic modeling task Byrne et al. (2022), and the model interpretation and analysis task Gourgoulias et al. (2024); Proskura and Zaytsev (2024). Gourgoulias et al. (2024) probes LLMs to estimate class separability of text datasets, and Proskura and Zaytsev (2024) uses topological information from encoder models to select the best models to create an ensemble. Finally, Rathore et al. (2023) performs the structure and visualization task by visualizing the training process of transformer-based models.
Insight (Transformer-CLS). CLS embedding is applied to four tasks - text classification (the most popular), topic modeling, model interpretation & analysis, and structure & visualization. The most interesting subtasks are probing LLMs to estimate class separability of text datasets, and visualizing the training process of transformer-based models.
@@
Hidden Weights. Classification tasks include deepfake text detection Uchendu et al. (2024), and LLM hallucination detection Bazarova et al. (2025). Next, researchers explore a combination of sentiment & semantic analysis and structure & visualization tasks by using Mapper to visualize polysemous words in the hidden representations of the BERT transformer model Garcia (2022). Similarly, Alexander and Wang (2023) combines the health analysis and visualization tasks to visualize GPT-3’s embeddings of hate speech, misinformation, and psychiatric disorder texts with Mapper. Also, Ruppik et al. (2024) performs the clustering and topic modeling task through dialogue term extraction. Finally, model interpretation and analysis is the most popular task, where Gardinazzi et al. (2024) proposes a novel metric - persistence similarity to prune redundant layers in LLMs; Balderas et al. (2025) proposes Persistent BERT Compression and Explainability (PBCE) to compress BERT by pruning redundant layers; Sun and Nelson (2023) probes the correlation between sentence vectors and their semantics; Chauhan and Kaul (2022) proposes a novel scoring metric - persistence scoring function which captures the homology of the hidden representations of BERT; García-Castellanos et al. (2024) performs zero-shot model stitching by employing topological densification; and Kudriashov et al. (2024) probes BERT’s hidden weights on new grammatical features, known as polypersonality.
Insight (Transformer-Hidden). Researchers have applied TDA features extracted from Hidden weights in to all tasks, except health, social, & scholarly analysis and speed & music processing tasks. From the various novel model interpretation and analysis applications, it seems that TDA features extracted from the Hidden weights are able to make these black-box models more transparent.
Attention Weights. BERT’s attention weights have been transformed to both directed and undirected graphs, on top of which different TDA features are extracted for text classification tasks such as deepfake text detection Kushnareva et al. (2021); Uchendu et al. (2024), robustness evaluation of TDA features Perez and Reinauer (2022), authorship attribution of Japanese texts Sakurai et al. (2025), out-of-distribution detection (OOD) Pollano et al. (2024); Perez and Reinauer (2022), and vulnerability detection in code Snopov and Golubinskiy (2024). Next, this framework is applied to the sentiment and semantic analysis task, specifically on human linguistic competence (i.e., grammatical acceptability judgment)  Cherniavskii et al. (2022); Proskurina et al. (2023); Perez and Reinauer (2022), dialog term extraction Vukovic et al. (2022), and document coherence Jain et al. (2024). Similarly, with the same framework, a speech and music processing application Tulchinskii et al. (2023). Finally, researchers attempt the model interpretation and analysis task, where Kostenok et al. (2023) uses the topological features extracted from the attention weights to estimate uncertainty in the encoder models, and Proskura and Zaytsev (2024) performs dynamic weighting for creating ensemble models.
Insight (Transformer-Attention). Given the wealth of information contained in the attention weights, researchers have applied topological features extracted from these weights to several tasks. The most interesting applications are observed on the sentiment and semantic tasks - grammatical acceptability judgment and document coherence analysis. Finally, we observe that for all the Transformer-based weights, the most explored application is deepfake text detection.
@@
ELMo Embeddings. ELMo embeddings are a type of word representation that captures both the meaning of words and their usage in context Peters et al. (2018). Similar to other embeddings, ELMo has also been leveraged to extract topological features. Tymochko et al. (2021) performs text classification by applying these embeddings for detecting fraudulent papers by examining their titles and abstracts. This is done in comparison of other embeddings (Word2Vec, GloVe, FastText, and Frequency Time Series) to ascertain the best embeddings to extract strong topological features.
Insight (ELMo). ELMo embeddings is applied to one task - text classification by one researcher in comparison to other embeddings.
@@ 4.2.4 Symbolic Representations
PLC refers to rules and methods used to encode letters that fuel various communication systems, cryptography techniques, or linguistic analyses. Letter coding transforms letters or characters into different symbols, numbers, or other forms. PSSC is similar to PLC but for extracting topological features from speech sounds. One particular application of PLC and PSSC is the study of tongue twisters, such as for Ukrainian Yurchuk and Gurnik (2023); Kovaliuk et al. (2024). These applications fall under the text classification task for PLC and speech & music processing task.
Insight (Symbolic). Symbolic representations are a novel technique for numerically representing text. The only applications of this technique are on Uranian tongue twisters, both for the classification and speech & processing tasks. This suggests that such a technique could be applied to processing low-resource languages, where numerical representation may be insufficient.
@@ 4.2.5 Multi-Modal Representations
TDA features have also been extracted from other representations of NLP-related features, including multimedia data such as audio and video. In this section, the only task performed by researchers is speech & music processing, where applications include - studying human vowels and infant vocalizations Bonafos et al. (2023, 2024), emotion recognition from audio speech Gonzalez-Diaz et al. (2019) and videos Paluzo-Hidalgo et al. (2022), depression detection from audio clips Tlachac et al. (2020), recognizing voiced and voiceless consonants in speech Zhu et al. (2024), music classification Bergomi (2015); Sassone et al. (2022), and assessing the adversarial robustness of image-text multi-modal models by measuring topological consistency Vu et al. (2025).
Insight (Multi-Modal). Multi-modal representations are another novel technique for numerically representing. The only task performed by researchers using this embedding is the speech & music processing task. This is because of the nature of the data being audio. Finally, the most interesting applications are the emotion recognition from audio speech, and depression detection from audio clips. All researchers show that topological features extracted from such embeddings experience high gains.
(a) Watermarked
(b) Non-Watermarked
Figure 6: Using Watermarked vs. Non-Watermarked Texts, (a) & (b) are the distance correlation matrix of TDA features vs. Linguistic features heatmaps
(a) Watermarked
(b) Non-Watermarked
Figure 7: Using Watermarked vs. Non-Watermarked Texts, (a) & (b) are Mapper plots to visualize the shape of data.
5 Case Studies of TDA for NLP Tasks
We showcase the utility of TDA in non-trivial NLP tasks by applying persistent homology and mapper using the Theoretical and Non-theoretical approaches. See below:
5.1 Theoretical Approaches.
(Semantic) Topological Space.
Although applications of this framework have yielded interesting findings, which are discussed in Section 3.1.1, it might still be interesting to answer the question - does topological features of text capture semantic features as well? Thus, we perform a small analysis to illustrate a real-world application. We perform this analysis on a relevant topic - distinguishing between watermarked and non-watermarked LLM-generated texts.
We extract topological and linguistic features for a correlation test. The text is initially represented numerically with BERT’s attention embeddings, then as an undirected graph, similar to Kushnareva et al. (2021) to extract topological features. Next, using the undirected graph as input, we calculate topological features in the 0-dimension, such as the number of non-zero values, maximum value, mean of values, number of values, and persistent entropy. Also, for the linguistic features, we extract entropy, perplexity, position of noun & adjective, coherence, flesh reading ease score & grade, number of stopwords, length of text, and number of unique words. Finally, we use distance correlation to measure the statistical dependence of the two non-linear sets of features. See Figures 6(a) & 6(b) for the correlation matrix heatmap. Many of the TDA features correlate with the linguistic features, with the highest correlation being the length of text (doc_length) and number of topological values in $\beta_{0}$ (betti_curve_0). Lastly, these results suggest that topological features can capture semantic features, however, a more comprehensive study is needed to make this a strong finding. In addition, such a finding can inform the non-theoretical approaches we adopt for real-world NLP tasks.
Topological “Shape” of Words.
Using the same method as above to represent the Watermarked vs. Non-Watermarked texts, we visualize the “shape” of their data using Mapper to observe the similarity or distinctness between the text types. See Figures 7(a) and 7(b) for the Mapper plots of the Watermarked and Non-Watermarked texts, respectively. We observe from the Figures that these different texts are very similar, although the watermarked text has slightly more nodes and edges than the non-watermarked texts, suggesting more connectivity. Thus, we can conclude that these texts are topologically similar, which could suggest that the watermarking technique does not significantly deviate from the style and quality of the non-watermarked version. However, to further explain these results, a more comprehensive study is required to interpret topological features extracted from text linguistically. This further stresses the validity of the theoretical study in Section 5.1. In addition, such findings could indicate that the watermarking technique does not break the semantics, style, and syntax of the non-watermarked pre-trained model.
5.2 Non-Theoretical Approaches
Distance Value (Avg $\pm$ std)
Euclidean 1.412 $\pm$ 0.0249
Wasserstein 0.054 $\pm$ 0.0204
Table 2: Average Euclidean and Wasserstein distances and standard deviation of the English-French text pairs
Model Interpretation and Analysis.
To illustrate an interesting task for applying TDA to NLP, we aim to investigate whether the embeddings of English and non-English text pairs are the same. This analysis can provide insights into the ability of text embeddings to capture the same information but in different languages. The idea is to represent an English text with a monolingual English embedding model and a non-English text with the same meaning with a multilingual embedding model, and compare the similarities between these embeddings. This similarity is captured by calculating both the Euclidean and topological distances. For the topological distance, we use Wasserstein distance, a similarity metric between two probability distributions, typically used in computational topology Panaretos and Zemel (2019); Shin (2019); Draganov and Skiena (2024). Specifically, we evaluate the Euclidean and Wasserstein distances between the embeddings of English-French translation text pairs. For the Euclidean distance, we compare the embeddings from the English model and the multilingual model. To calculate the Wasserstein distance, we first extract topological features from the embeddings, represent them as persistent diagrams, and then calculate the distance. Euclidean and Wasserstein distance range is [0, $\infty$). We further normalize the embeddings such that both distance metrics are comparable. See Table 2 for the distance values. Thus, we obtain an average distance of 1.412 and 0.054 for the Euclidean and Wasserstein distances, respectively. These values suggest that these embeddings capture the similarity between the text pairs, with Wasserstein being almost zero, meaning that the text pair embeddings are topologically similar. Finally, we propose a new way of evaluating the strength of multilingual models by observing their topological similarity to monolingual models.
Pros Cons
Captures Global & Local Text Structure – TDA models relationships between words, sentences, or documents in a topological space, uncovering hierarchical and contextual structures. High Computational Cost – Applying TDA to large text corpora requires significant processing power, especially when constructing high-dimensional topological features.
Robust to Noise & Variability – Persistent homology can filter out minor textual variations while retaining essential linguistic structures. Interpretability Challenges – Persistence diagrams and barcodes are not intuitive for NLP practitioners, requiring additional processing to extract meaningful linguistic insights.
Effective for Low-Resource Settings – Unlike deep learning, which requires large datasets, TDA can work with smaller corpora by leveraging topological structures rather than statistical frequency-based methods. Limited Software & NLP-Specific Tools – Most TDA tools are designed for point clouds and image data, requiring adaptations for NLP applications.
Detects Complex Relationships – TDA can uncover semantic relationships and linguistic patterns that traditional word embeddings may miss. Not Yet Standard in NLP Pipelines – TDA is still experimental in NLP, lacking standardized frameworks for integration with common NLP libraries like spaCy, Transformers, or NLTK.
No Need for Explicit Feature Engineering – Unlike many traditional methods, TDA analyzes raw data without requiring predefined features. Limited Adoption – TDA is still an emerging field, meaning fewer case studies and industrial applications in NLP compared to traditional methods.
Works Well with Word Embeddings & Transformers – Can be integrated with word2vec, and Encoder & Decoder model embeddings to enhance understanding of text structures and improve classification tasks. Requires Specialized Expertise – Implementing TDA in NLP requires both topology and NLP expertise, making adoption difficult for standard NLP practitioners.
Provides Geometric Insights into Syntax & Semantics – Helps visualize the shape of linguistic structures, which is valuable in understanding complex texts. Difficult to Benchmark – Unlike traditional NLP metrics (e.g., BLEU, perplexity), there is no clear evaluation standard for TDA-based NLP models.
Table 3: Pros and Cons of Topological Data Analysis (TDA) in NLP